OpenStudy (anonymous):

3/x +(x+1/x+2) Simplify

OpenStudy (anonymous):

\[\frac{ 3 }{x } + \frac{ x+1 }{ x+2 }\]

OpenStudy (tkhunny):

You have written \(\dfrac{3}{x} + x + \dfrac{1}{x} + 2\). Was this your intent? Perhaps you meant 3/x + (x+1)/(x+2) = \(\dfrac{3}{x} + \dfrac{x+1}{x+2}\) Do you see that your original notation is inadequate?

OpenStudy (tkhunny):

I suggest you find a common denominator and add the two fractions.

OpenStudy (anonymous):

How do you do that though?

OpenStudy (tkhunny):

Have you considered x(x+2) as the common denominator?

OpenStudy (anonymous):

I have no Idea how to do this, could you go through it step by step?

OpenStudy (anonymous):

oh, ok... that kinda helped but that wasn't a good explanation

OpenStudy (tkhunny):

You cannot judge the value of an explanation. You can judge only what you get and what you don't. Please add these two fractions: \(\dfrac{1}{3} + \dfrac{1}{5}\)

OpenStudy (anonymous):

\[\frac{ 8 }{ 15 }\]

OpenStudy (tkhunny):

Do this one EXACTLY the same. Find the common denominator and multiply each term by whatever it is missing in the denominator. EXACTLY the same idea.